The average price of Yankee season ticket behind home plate was 27,858 in 2011. nine years earlier, the average price was 21308. What was the annual increase in price?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

27,858=21308*(1+r/100)^9

(27,858/21308)^(1/9)=(1+r/100)

(1+r/100)=1.0302

r=1.0302-1

=3.02%(Approx).